OpenStudy (anonymous):

In the following reaction, how many grams of NaBr will react with 311 grams of Pb(NO3)2? Pb(NO3)2(aq) + 2NaBr(g) d PbBr2(s) + 2NaNO3(aq) The molar mass of NaBr is 102.9 grams and that of Pb(NO3)2 is 331.21 grams.

OpenStudy (anonymous):

The steps to follow are balance the equation and recognise the molar ratios or mass ratios. For this problem, let us work with mass ratios. As the equation suggests, the molar ratio is 1:2 of Pb(NO3)2 to NaBr and in terms of masses shall be 331.21 : (2*102.9) or 331.21 : 205.8 Simply, if we are given 311 of the lead compoud, then that of NaBr= (311*331.21)/205.8= 500.5 grams.

OpenStudy (jfraser):

@hoslos mass ratios are \(terrible\) for use in chemistry, we need to use the molar ratios. Your mass makes no sense, because your mass ratio is upside down. You wouldn't know it because you're missing the units that go along with the conversion

OpenStudy (anonymous):

The masses are all given in grams. Besides, I have been using this technique to solve IGCSE chemistry papers where necessary and they all worked out.

OpenStudy (jfraser):

\[311g Pb(NO_3)_2 * (\frac{1mol Pb(NO_3)_2}{331.2g})*(\frac{2mol NaBr}{1mol Pb(NO_3)_2})*(\frac{102.9g NaBr}{1mol NaBr})\] is 193g, not 500g

OpenStudy (anonymous):

Your method is correct. Mine is too. I just exchanged some numbers. It sould have been (311*205.8)/331.21= 193.2grams. Hence it is the same, not be called TERRIBLE.

OpenStudy (jfraser):

using mass ratios isn't good chemistry

OpenStudy (anonymous):

But I told they have worked out . You simply apply the same rule as of ratio. Wether if it 1:2 or what. Plus, the same applies to volumes, not only masses.